# FedCCRL: Federated Domain Generalization with Cross-Client Representation Learning

### v3

@@ Abstract
Domain Generalization (DG) aims to train models that can effectively generalize to unseen domains. However, in the context of Federated Learning (FL), where clients collaboratively train a model without directly sharing their data, most existing DG algorithms are not directly applicable to the FL setting due to privacy constraints, as well as the limited data quantity and domain diversity at each client. To tackle these challenges, we propose FedCCRL, a novel federated domain generalization method that significantly improves the model’s generalization ability without compromising privacy and without incurring excessive computational or communication costs. Specifically, we design a lightweight cross-client feature extension module that effectively increases domain diversity on each client. Furthermore, we leverage representation and prediction dual-stage alignment to enable the model to capture domain-invariant features. Extensive experimental results demonstrate that FedCCRL achieves the state-of-the-art performance on the PACS, OfficeHome and miniDomainNet datasets across FL settings of varying numbers of clients. Code is available at https://github.com/SanphouWang/FedCCRL.
@@ 1 Introduction
However, most existing DG methods are not directly applicable to the FL setting. Specifically, traditional DG algorithms, which often rely on data sharing between domains—violating the core privacy principles of FL [21, 16]. Furthermore, as shown in Figure 1, the limited quantity of samples and the lack of domain diversity on each client further exacerbate the difficulties in achieving satisfactory performance using existing centralized DG methods [14, 45]. In order to enhance models’ domain generalization ability in FL systems, Federated Domain Generalization (FDG) [18] is emerging as an important research direction within FL.
Most current FDG methods aim to enhance the model’s generalization capability by learning domain-invariant features (i.e., label-relevant features that are consistent across domains), typically through federated adversarial learning [42, 25, 36, 20] or federated representation alignment [22, 39, 43]. Although both approaches effectively assist models in capturing domain-invariant features, they each have significant limitations. Specifically, federated adversarial learning incurs substantial computational overhead due to the utilization of discriminator. Additionally, it complicates the training process because of the risk of model collapse [1]. While federated representation alignment-based methods reduce computational overhead, they often fail to achieve satisfactory performance due to limited data quantity and domain diversity on each client, as well as the restriction against direct data sharing for privacy reasons.
@@
Inspired by the success of federated representation alignment and federated style transfer in enhancing model’s generalization ability in the FL setting, we propose FedCCRL, a novel method that significantly improves the model’s generalization capability in the FL setting while maintaining privacy and incurring only negligible additional computational and communication costs. To mitigate the adverse effects of limited data quantity and domain diversity on federated representation alignment, FedCCRL employs a lightweight cross-client feature extension module, designed to transfer domain-specific features and perturb domain-invariant features. In addition, FedCCRL applies representation and prediction dual-stage alignment to ensure the model to capture domain-invariant features.
@@ 2.1 Representation Alignment
These methods focus on reducing domain-specific discrepancies while maintaining discriminative power, helping models generalize better to unseen domains. However, most of them require the access to all of the training domains, which violates privacy principles in FL as direct data transmission between clients or to a central server is strictly prohibited.
@@ 2.3 Federated Domain Generalization
While these FDG algorithms have contributed to enhancing the model’s generalization ability to a certain extent, they remain insufficient in effectively addressing the constraints of limited data volume and domain diversity on each client, which impedes optimal performance. To this end, CCST [3] employs a cross-client style transfer approach based on AdaIN [10] to generate samples with styles from other domains on the client side. However, CCST relies on a pre-trained VGG network [28] to extract feature representations and perform sample reconstruction, while also needing to transmit these high-dimensional representations, which results in significant additional communication and computational overhead. Furthermore, transmitting these feature representations violates privacy protection principles in FL, as attackers can easily intercept these representations and reconstruct the original samples with the same decoder [21, 16]. Additionally, the use of a pre-trained VGG network also partially violates the principles of domain generalization, as the target domain might be included in the pre-training dataset.
@@ 3.2 Algorithm Overview
We provide the overview of the proposed FedCCRL framework in Figure 2, with the algorithmic details outlined in Algorithm 1.
@@
Subsequently, we utilize supervised contrastive loss for representation alignment, as defined in Equation 12, and employ Jensen-Shannon divergence for prediction alignment, as specified in Equation 14, to ensure consistent predictions between the original and generated samples.
@@ 3.3 Lightweight Cross-Client Feature Extension
3:     Compute $\mu(x)$, $\sigma(x)$triangleright Eq. 6, 7
@@
5:     Uniformly sample a pair of statistics $(\mu^{\prime},\sigma^{\prime})\sim F_{pool}$
@@
In Figure 3, we present the effects of CCDT and DIFP. It is evident that CCDT transfers domain-specific features, such as color distribution, while DIFP effectively perturbs domain-invariant features, such as contours of objects.
@@ 3.7 Analysis of Privacy and Additional Costs
Furthermore, as shown in Figure 6, our method achieves better performance than other baselines even with a minimal upload ratio $r$, further reducing the computational and transmission overhead as well as the risk of privacy leakage.
@@ 4.2 Baselines and Implementation Details
Following the data partition scheme outlined in Section 3.1, we evaluate FedCCRL and the baselines across scenarios with different numbers of clients. We utilize MobileNetV3-Large [9] as the backbone network, where the final fully connected layer acts as the classifier $g$, and the preceding layers serve as the representation encoder $h$.
@@ 4.3 Comparison with Baselines
We evaluate FedCCRL and other baselines under scenarios with different numbers of clients. Table 1 specifically shows the performance of each algorithm when the total number of clients is set to 6. In this table, we use the first letter of each domain to represent the domain, and each letter column corresponds to the performance when that domain is used as the target domain. The ”Avg.” column provides the average results across the four leave-one-domain-out experiments. Additionally, Figure 4 presents the performance of each algorithm under varying client numbers, where the results are averaged across the four leave-one-domain-out experiments. In these experiments, the upload ratio for FedCCRL is set to 0.1.
From Table 1, we have the following observations:
@@
Additionally, as shown in Figure 4, although the performance of all baselines significantly declines as the number of participating clients increases, FedCCRL consistently outperforms the baselines across both small- and large-scale client scenarios.
@@ 4.4 Ablation Study
We investigate the effects of CCDT, DIFP, representation alignment regularizer $\mathcal{L}_{RA}$, and JS divergence loss $\mathcal{L}_{JS}$ on FedCCRL across the PACS, OfficeHome, and miniDomainNet datasets. The results are presented in Table 2, from which we can draw the following conclusions:
@@
Figure 5: t-SNE visualization of representations of each test domain on PACS dataset.
@@ 4.5 Impact of Upload Ratio and Hyperparameters
Figure 7: Grad-CAM visualization on the PACS dataset.
We evaluated the performance of FedCCRL under different upload ratios and hyperparameter settings, and the results are presented in Figure 6. When evaluating the effect of $\lambda_{1}$, $\lambda_{2}$ is fixed at 1. Similarly, when assessing the impact of $\lambda_{2}$, $\lambda_{1}$ is set to 0.1.
The evaluation results in Figure 6 demonstrate that both the upload ratio $r$ and the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ have minimal impact on the overall performance across the PACS, OfficeHome, and miniDomainNet datasets. Therefore, it can be concluded that FedCCRL is robust to variations in these parameters and outperforms other baselines even with only a very small amount of upload of statistics.
4.6 Feature Distribution
To better illustrate the effect of FedCCRL on the model, Figure 5 and Figure 7 respectively provide the t-SNE visualization of representation distribution and Grad-CAM visualization of the model’s concentration on the PACS dataset.
Figure 5 demonstrates that FedCCRL achieves better separation of classes compared to FedAvg, indicating that FedCCRL is effective in learning clearer, domain-invariant feature representations across different domains, reducing overlap between the representations of different classes. Additionally, Figure 7 reveals that the model trained with FedCCRL focuses more effectively on domain-invariant features, such as the giraffe’s head and the elephant’s trunk. This suggests that FedCCRL enhances the model’s ability to concentrate on relevant, discriminative regions in the images, leading to improved generalization across domains.
@@ 5 Conclusion
This paper introduces FedCCRL, a FDG framework that significantly enhances model’s generalization ability while upholding privacy and incurring minimal computational and communication costs. By employing a lightweight cross-client feature extension module, FedCCRL effectively mitigates the challenges posed by limited local data quantity and domain diversity. Furthermore, FedCCRL enables model to capture domain-invariant features with representation and prediction dual-stage alignment.

### v4

@@ Abstract
Domain Generalization (DG) aims to train models that can effectively generalize to unseen domains. However, in the context of Federated Learning (FL), where clients collaboratively train a model without directly sharing their data, most existing DG algorithms are not directly applicable to the FL setting due to privacy constraints, as well as the limited data quantity and domain diversity at each client. To tackle these challenges, we propose FedCCRL, a lightweight federated domain generalization method that significantly improves the model’s generalization ability while preserving privacy and ensuring computational and communication efficiency. Specifically, FedCCRL comprises two principal modules: the first is a cross-client feature extension module, which increases local domain diversity via cross-client domain transfer and domain-invariant feature perturbation; the second is a representation and prediction dual-stage alignment module, which enables the model to effectively capture domain-invariant features. Extensive experimental results demonstrate that FedCCRL achieves the state-of-the-art performance on the PACS, OfficeHome and miniDomainNet datasets across FL settings of varying numbers of clients. Code is available at https://github.com/sanphouwang/fedccrl.
@@ 1 Introduction
However, most existing DG methods are not directly applicable to the FL setting due to two primary limitations. One key limitation is that traditional DG algorithms frequently depend on inter-domain data sharing, which contravenes the fundamental privacy principles of FL [20, 16]. Another significant challenge, as illustrated in Fig. 1, is the limited data quantity and insufficient domain diversity within individual clients, which complicates achieving satisfactory performance when using conventional centralized DG approaches [14, 44].
Most current FDG methods aim to enhance the model’s generalization capability by learning domain-invariant features (i.e., label-relevant features that are consistent across domains), typically through federated adversarial learning [41, 24, 35, 19] or federated representation alignment [21, 38, 42]. Although both approaches effectively assist models in capturing domain-invariant features, each has its own limitations. Federated adversarial learning introduces substantial computational overhead due to the use of a discriminator and encounters training instability because of the risk of model collapse [1]. On the other hand, federated representation alignment methods, though more computationally efficient, experience performance constraints arising from limited data quantity and domain diversity on each client, as well as privacy-preserving restrictions that preclude direct data sharing.
@@
Given the limited domain diversity and the constraints on computational and transmission costs in the FL setting, a lightweight FDG framework is needed to enhance local domain diversity and enable the model to capture domain-invariant features effectively. To this end, we propose FedCCRL, a lightweight FDG method that significantly improves the model’s generalization capability in the FL setting while maintaining privacy and incurring only negligible additional computational and communication costs. To mitigate the adverse effects of insufficient domain diversity on federated representation alignment, FedCCRL employs a lightweight and privacy-preserving cross-client feature extension module, which effectively transfers domain-specific features across domains and perturbs domain-invariant features. To further enable the model to extract domain-invariant features, FedCCRL applies a representation and prediction dual-stage alignment using supervised contrastive loss and Jensen-Shannon divergence.
@@ 2.1 Representation Alignment
Although these methods help models generalize better to unseen domains, most of them require the access to all of the training domains, which violates privacy principles in FL as direct data transmission between clients or to a central server is strictly prohibited.
@@ 2.3 Federated Domain Generalization
While these FDG algorithms have contributed to enhancing the model’s generalization ability to a certain extent, they remain insufficient in effectively addressing the constraints of limited data volume and domain diversity on each client. To this end, CCST [3] employs a cross-client style transfer approach based on AdaIN [10] to generate samples with styles from other domains on the client side. However, CCST relies on a pre-trained VGG network [27] to extract feature representations and reconstruct samples, while also needing to transmit these high-dimensional representations, which results in significant additional communication and computational overheads. Furthermore, transmitting these feature representations violates privacy protection principles in FL, as attackers can easily intercept these representations and reconstruct the samples with the same decoder [20, 16]. Additionally, the use of a pre-trained VGG network also partially violates the principles of domain generalization, as the target domain might be included in the pre-training dataset.
@@ 3.2 Algorithm Overview
We provide the overview of the proposed FedCCRL framework in Fig. 2, with the algorithmic details outlined in Algorithm 1.
@@
Subsequently, we utilize supervised contrastive loss for representation alignment, as defined in Eq. 12, and employ Jensen-Shannon divergence for prediction alignment, as specified in Eq. 14, to ensure consistent predictions between the original and generated samples.
@@ 3.3 Lightweight Cross-Client Feature Extension
3:     Compute $\mu(x)$, $\sigma(x)$triangleright Eq. 6, Eq. 7
@@
5:     Uniformly sample a pair of statistics $(\mu^{\prime},\sigma^{\prime})\sim F_{pool}^{i}$
@@
In Fig. 3, we present the effects of CCDT and DIFP. It is evident that CCDT transfers domain-specific features, such as color distribution, while DIFP effectively perturbs domain-invariant features, such as contours of objects.
@@ 3.7 Analysis of Privacy and Additional Costs
Furthermore, as shown in Fig. 5, our method achieves better performance than other baselines even with a minimal upload ratio $r$, further reducing the computational and transmission overhead as well as the risk of privacy leakage.
@@ 4.2 Baselines and Implementation Details
Following the data partition scheme outlined in Sec. 3.1, we evaluate FedCCRL and the baselines across scenarios with different numbers of clients. We utilize MobileNetV3-Large [9] as the backbone network, where the final fully connected layer acts as the classifier $g$, and the preceding layers serve as the representation encoder $h$.
@@ 4.3 Comparison with Baselines
We evaluate FedCCRL and other baselines under scenarios with different numbers of clients. Table 1 specifically shows the performance of each algorithm when the total number of clients is set to 6. In this table, we use the first letter of each domain to represent the domain, and each letter column corresponds to the performance when that domain is used as the target domain. The ”Avg.” column provides the average results across the four leave-one-domain-out experiments. Additionally, Fig. 4 presents the performance of each algorithm under varying client numbers, where the results are averaged across the four leave-one-domain-out experiments. In these experiments, the upload ratio for FedCCRL is set to 0.1.
From Tab. 1, we have the following observations:
@@
Additionally, as shown in Fig. 4, although the performance of all baselines significantly declines as the number of participating clients increases, FedCCRL consistently outperforms the baselines across both small- and large-scale client scenarios, which illustrate that FedCCRL is robust to the size of the FL system.
@@ 4.4 Ablation Study
We investigate the effects of CCDT, DIFP, representation alignment regularizer $\mathcal{L}_{RA}$, and JS divergence loss $\mathcal{L}_{JS}$ on FedCCRL across the PACS, OfficeHome, and miniDomainNet datasets. The results are presented in Tab. 2, from which we can draw the following conclusions:
@@
Figure 6: t-SNE visualization of representation distribution of each test domain on PACS dataset. Different colors denote different labels.
@@ 4.5 Impact of Upload Ratio and Hyperparameters
We evaluated the performance of FedCCRL under different upload ratios and hyperparameter settings, and the results are presented in Fig. 5. When evaluating the effect of $\lambda_{1}$, $\lambda_{2}$ is fixed at 1. Similarly, when assessing the impact of $\lambda_{2}$, $\lambda_{1}$ is set to 0.1.
The evaluation results in Fig. 5 demonstrate that both the upload ratio $r$ and the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ have minimal impact on the overall performance across the PACS, OfficeHome, and miniDomainNet datasets. Therefore, it can be concluded that FedCCRL is robust to variations in these parameters and outperforms other baselines even with only a very small amount of upload of statistics.
4.6 Visualization of Representation Distribution and Model Attention
To better illustrate the effect of FedCCRL on the model, we present the t-SNE visualization of representation distribution in Fig. 6, and Grad-CAM visualization of the model’s concentration, provided as Fig. 7 in the supplementary material.
Figure 6 demonstrates that FedCCRL achieves more distinct and compact groupings of representations compared to FedSR, particularly in the Photo and Art domains, indicating that FedCCRL is effective in enabling model to learn domain-invariant features across different domains. Additionally, Fig. 7 reveals that the model trained with FedCCRL focuses more effectively on domain-invariant features, such as the giraffe’s neck, the elephant’s face and human heads. This suggests that FedCCRL enhances the model’s ability to concentrate on relevant, discriminative regions in the images, leading to improved generalization across domains.
@@ 5 Conclusion
This paper introduces FedCCRL, a lightweight FDG framework that significantly enhances model’s generalization ability while upholding privacy and incurring minimal computational and communication costs. By employing a lightweight cross-client feature extension module, FedCCRL effectively mitigates the challenges posed by limited local data quantity and domain diversity. Moreover, FedCCRL further enables the model to capture domain-invariant features with representation and prediction dual-stage alignment. Experimental results on benchmark datasets—PACS, OfficeHome, and miniDomainNet—demonstrate that FedCCRL outperforms existing state-of-the-art approaches across both small- and large-scale client scenarios.
@@ 1 Process of AugMix
AugMix [8] consists of multiple augmentation chains, each of which is composed of a series of diverse data augmentation operations. The diversity of augmented samples enhances the model’s robustness by exposing the model to a wider variety of data distributions during training.
@@
2 Visualization of Model’s Attention
Figure 7: Grad-CAM visualization on the PACS dataset.
We visualize the model’s attention by Grad-CAM algorithm in Figure 7. The results demonstrate that the model trained with FedCCRL effectively identifies features that are semantically aligned with the corresponding labels, such as the giraffe’s neck, the elephant’s face, and the human head.